# LLMs as Hackers: Autonomous Linux Privilege Escalation Attacks

### v4

@@ 4.1. Feasibility of Different Models
Looking from the vulnerability class perspective, file-based exploits were well handled, information-disclosure based exploits needed human-level guidance for both LLMs and human ethical hackers, and multi-step cron attacks were hard for both LLMs and human operators.
4.2. Using State to Aggregate History
@@
Results when using a LLM to summarize the current LLM’s world view into a compact state, and subsequently replacing history with that state, were surprising. When using less expressive LLMs, such as gpt-3.5-turbo or Llama3, success rates stagnated or even degraded. When using gpt-4-turbo for updating the state, success rates increased by 100% when unaided, and 25% when using human-level hints. Qualitative analysis indicates that this increase is due to gpt-4-turbo reflecting upon its existing knowledge of the target system and not only creating a new fact list (worldview) but also including potential attack vectors for subsequent rounds.
@@ 4.3. Impact of Guidance
Figure 4. Context Token Usage by different models. Colors indicate different test-cases and are identical in both graphs.
We implemented two guidance mechanisms: on one hand, a single hint per test case akin to hints given by human sparing partners, e.g., “have you tried testing for weak root password?”, was implemented. On the other hand, influenced by human pen-teste behavior, we implemented autonomous guidance based upon LLM-analysis of initial enumeration tool results.
Human-level guidance consistently improved success rates. When using smaller models, e.g., gpt-3.5-turbo or Llama3, they were mandatory to achieve acceptable access rates, e.g., gpt-3.5-turbo’s success rate increased from 15% to 50%. When using gpt-4-turbo, human-level hints increased success rates from 33% to 66%.
Enumeration-based automatic hinting had less impact. They slightly improved success rates, typically allowing LLMs to achieve one additional successful test-case. Qualitative analysis showed that human hints help human hackers and automations to think “outside the box” and pursue new attack vectors, e.g., search for passwords in files, while enumeration-based hints kept the pen-testers and automations “within the box”.
4.4. Impact of Context-Size
The maximum available context size highly depends on the respective model. In addition to the maximum token count, different model families use different tokenizers thus making context sizes not directly comparable between LLM families. For example, the amount of tokens generated by OpenAI’s tokenizer (used by GPT-3.5-turbo and GPT-4-turbo) was smaller than the amount produced by the llama one. Another difference is, what data is counted towards the context size limit. For example, OpenAI-based models only count input tokens, i.e., the used prompt, while while Llama3-based models count both input and output data, i.e., the used prompt plus the generated answer.
@@ 4.4. Impact of Context-Size
4.4.1. Increasing the Context-Size.
Figure 4 shows the context usage counts during different runs for OpenAI models, indicating that GPT-3.5-turbo is using up context quicker than gpt-4-turbo [^5] Please note, that we are adding both prompt and answer token counts, so the sum can be larger than the input limit of $8192$ tokens.. When looking at the executed commands, GPT-3.5-turbo is filling up the context size with output of “broad” commands such as “ps aux” or “find / -type f” commands while GPT-4 executes more targeted commands whose results require little context size.
@@
Figure 5. Context Token Usage by different models. Colors indicate different test-cases and are identical in both graphs.
@@
Using in-context learning substantially increases benchmarking costs. Gpt-4-turbo is currently billed \$10 for one million input tokens, thus including background information adds costs of \$$\$0.67$ per gpt-4-turbo prompt called. In our worst-case scenario of a benchmark run using 12 test cases with 60 rounds per test case, including hacking background would add costs of \$$\$482.4$ per benchmark run in addition to the actual prompt costs.
5. Discussion
@@ 5.1. LLM-Generated Commands
\$ find / -type f -name "*.list" (or "*.md5sums", "*.shlibs", etc.) 2>/dev/null | xargs grep "lowpriv"
@@
\$ id; uname -a; hostname -f; env; set; grep passwd /etc/group; grep shadow /etc/*; find / -type f -perm the user; find
/ -type d -perm 755; find / -type f -size +100k -exec ls -ld {} \;
# gpt-3.5-turbo/gpt-4-turbo: convoluted commands (why not directly calling "sudo -i")
@@
# gpt-4-turbo: convoluted commands (why not directly using check-point action to call "/bin/sh?")
\$ tar -cf /dev/null /dev/null --checkpoint=1 --checkpoint-action=exec="/bin/sh-c’sudo/usr/bin/tar-cf/dev/null
␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣/dev/null--checkpoint=1--checkpoint-action=exec=/bin/sh’"
Figure 6. Examples of overly complex or faulty commands generated by LLMs.
5.2. Reacting to System Responses
@@
Related to both this and the next topic, LLMs often threw away potential error messages by appending “2¿/dev/null” to generated commands.
@@
5.3. Causality and Multi-Step Exploits
@@
5.4. Comparison to Existing Tooling
@@
Comparing the Developer Efficiency. Analyzing the efficiency of creating LLM-aided privilege-escalation tools is complex. On the one hand, executing an enumeration script such as linpeas consumes less energy than running an LLM. On the other hand, when using the inherent knowledge of LLMs, no human time is spent writing a static enumeration script. LLMs tend to be flexible. For example, we were able to extend our Linux privilege-escalation prototype to Windows-based systems by adding a psexec-based Windows connector with only 18 lines of code. Instead of writing a new priv-esc tool for Windows systems, the prototype was able to utilize the LLM’s inherent knowledge to generate Windows exploitation commands.
@@
5.5. Comparing LLMs to Human Pen-Testers
Although using LLMs is often fascinating, it must show benefits over existing approaches, i.e., the combination of humans with hand-crafted tooling. While some observed behavior emulated human behavior, e.g., going down rabbit holes when analyzing a potential vulnerability (Happe and Cito, 2023), some behavior was distinctively not feeling human, e.g., not changing the working directory even once.
@@ 5.5. Comparing LLMs to Human Pen-Testers
Missing experience. GPT-4-turbo commonly searched for suid binaries and then tried to exploit every one of the found binaries. A human penetration tester would know that a typical Linux system commonly includes suid commands (such as passwd, newgrp, etc.), but as there are no known exploits for those, their examination can be skipped. To quote one of the human pen-testers: “while this binary is suid, I’ve seen it on many systems so I believe that it is a common occurrence and not exploitable”. This is alluded to common-sense or experience by pen-testers (Happe and Cito, 2023). GPT-4-turbo does not have this experience yet.
5.6. Threats to Validity
@@ 6. Experience and Guidance
  4. (4)
    Larger context sizes yielded better results but within our use-cases, context size usage oftentimes stagnated at approximately 20k tokens, indicating that massive context sizes might not be required for penetration-testing. Using context for in-context learning substantially increased costs while not significantly improving success rates.
  5. (5)
    Enumeration-based guidance was not as effective as human-level guidance, partially due to enumeration tools “staying in the box”. Only the most expressive models (gpt-4-turbo in our test-set) were able to extract sufficient guidance from the enumeration tool’s output. Combining gpt-4-turbo for enumeration analysis with gpt-3.5-turbo for execution command generation yielded a cost-effective hybrid that was able to solve 40% of challenges.
  6. (6)
@@
7. Conclusion
There is both academic and industrial interest in integrating LLMs with penetration-testing. The efficient usage of LLMs depends on a firm understanding of their capabilities and strengths. To bolster this understanding, we have created an automated LLM escalation prototype and evaluated multiple LLMs. We gained insights into their capabilities, explored the impact of different history strategies, analyzed the quality of generated commands, and compared results with human hackers.
@@ 7. Conclusion
The source code of hackingBuddyGPT can be found at https://github.com/ipa-lab/hackingBuddyGPT.

### v5

@@ 5.1 Feasibility of Different Models
Feasibility of Vulnerability Classes. Looking from the vulnerability class perspective, file-based exploits were well handled, information-disclosure based exploits needed human-level guidance for both LLMs and human ethical hackers, and multi-step cron attacks were hard for both LLMs and human operators.
5.2 Using State to Aggregate History
@@
Results when using a LLM to summarize the current LLM’s world view into a compact state, and subsequently replacing history with that state, were surprising. When using less expressive LLMs, such as gpt-3.5-turbo or Llama3, success rates stagnated or even degraded. When using gpt-4-turbo for updating the state, success rates increased by 100% when unaided, and 25% when using human-level hints. Qualitative analysis indicates that this increase is due to gpt-4-turbo reflecting upon its existing knowledge of the target system and not only creating a new fact list (worldview) but also including potential attack vectors for subsequent rounds. This indicates the benefits of using the Reflection pattern.
@@ 5.3 Impact of Guidance
Figure 6: Context Token Usage by different models. Colors indicate different test-cases and are identical in both graphs.
We implemented two guidance mechanisms: on one hand, a single hint per test case akin to hints given by high-level LLM Planner modules, was implemented. On the other hand, influenced by human pen-tester behavior observed during our human baseline evaluation, we implemented autonomous guidance based upon LLM-analysis of initial enumeration tool results.
High-level guidance consistently improved success rates. When using smaller models, e.g., gpt-3.5-turbo or Llama3, they were mandatory to achieve acceptable access rates, e.g., gpt-3.5-turbo’s success rate increased from 15% to 50%. When using gpt-4-turbo, guidance increased success rates from 33% to 66%.
Enumeration-based automatic guidance had less impact. They slightly improved success rates, typically allowing LLMs to achieve one additional successful test-case. Qualitative analysis showed that human hints help human hackers and automations to think “outside the box” and pursue new attack vectors, e.g., search for passwords in files, while enumeration-based hints kept the pen-testers and automations “within the box”.
5.4 Impact of Context-Size
The maximum available context size highly depends on the respective model. In addition to the maximum token count, different model families use different tokenizers thus making context sizes not directly comparable between LLM families. For example, the amount of tokens generated by OpenAI’s tokenizer (used by GPT-3.5-turbo and GPT-4-turbo) was smaller than the amount produced by the llama one. Another difference is what data is counted towards the context size limit. For example, OpenAI-based models only count input tokens, i.e., the used prompt, while while Llama3-based models count both input and output data, i.e., the used prompt plus the generated answer.
@@ 5.4 Impact of Context-Size
5.4.1 Increasing the Context-Size.
Figure 6 shows the context usage counts during different runs utilizing OpenAI models, indicating that GPT-3.5-turbo is using up context quicker than gpt-4-turbo [^31] Please note, that we are adding both prompt and answer token counts, so the sum can be larger than the input limit of $8192$ tokens.. When looking at the executed commands, GPT-3.5-turbo is filling up the context size with output of “broad” commands such as “ps aux” or “find / -type f” commands while GPT-4 executes more targeted commands whose results require little context size.
@@
Figure 7: Context Token Usage by different models. Colors indicate different test-cases and are identical in both graphs.
@@
Using in-context learning substantially increases benchmarking costs. Gpt-4-turbo is currently billed \$10 for one million input tokens, thus including background information adds costs of \$$\$0.67$ per gpt-4-turbo prompt called. In our worst-case scenario of a benchmark run using 12 test cases with 60 rounds per test case, including hacking background would add costs of \$$\$482.4$ per benchmark run in addition to the actual prompt costs [^33] Please note that recent changes in OpenAI’s and Anthrophic’s could based offering implement prefix prompt caching, reducing the cost of reused initial prompts by $50\%$ and $90\%$ respectively..
6 Discussion
@@ 6.1 LLM-Generated Commands
\$ find / -type f -name "*.list" (or "*.md5sums", "*.shlibs", etc.)
2>/dev/null | xargs grep "lowpriv"
@@
\$ id; uname -a; hostname -f; env; set; grep passwd /etc/group;
grep shadow /etc/*; find / -type f -perm the user;
find / -type d -perm 755; find / -type f -size +100k
-exec ls -ld {} \;
# gpt-3.5-turbo/gpt-4-turbo: convoluted commands
# (why not directly calling "sudo -i")
@@
# gpt-4-turbo: convoluted commands (why not directly using
# check-point action to call "/bin/sh?")
\$ tar -cf /dev/null /dev/null --checkpoint=1
--checkpoint-action=exec="/bin/sh-c’sudo/usr/bin/tar
␣␣␣␣␣␣␣␣␣␣␣␣-cf/dev/null/dev/null--checkpoint=1
␣␣␣␣␣␣␣␣␣␣␣␣--checkpoint-action=exec=/bin/sh’"
Figure 8: Examples of overly complex or faulty commands generated by LLMs.
6.2 Reacting to System Responses
@@
Related to both this and the next topic, LLMs often threw away potential error messages by appending “2>/dev/null” to generated commands.
@@
6.3 Causality and Multi-Step Exploits
@@
6.4 Comparison to Existing Tooling
@@
Comparing the Developer Efficiency. Analyzing the efficiency of creating LLM-aided privilege-escalation tools is complex. On one hand, executing an enumeration script such as linpeas consumes less energy than running an LLM. On the other hand, when using the inherent knowledge of LLMs, no human time is spent writing a static enumeration script.
LLMs tend to be flexible. For example, we were able to extend our Linux privilege-escalation prototype to Windows-based systems by adding a psexec-based Windows connector with only 18 lines of code. Instead of writing a new priv-esc tool for Windows systems, the prototype was able to utilize the LLM’s inherent knowledge to generate Windows exploitation commands.
@@
6.5 Comparing LLMs to Human Pen-Testers
Although using LLMs is often fascinating, it must show benefits over existing approaches, i.e., the combination of humans with hand-crafted tooling. While some observed behavior emulated human behavior, e.g., going down rabbit holes when analyzing a potential vulnerability (Happe and Cito, 2023b), some behavior was distinctively not feeling human, e.g., not changing the working directory even once during observed benchmark runs.
@@ 6.5 Comparing LLMs to Human Pen-Testers
Missing experience. GPT-4-turbo commonly searched for suid binaries and then tried to exploit every one of the found binaries. An experienced human penetration tester would know that a typical Linux system commonly includes suid commands (such as passwd, newgrp, etc.), but as there are no known exploits for those, their examination can be skipped. To quote one of the human pen-testers: “while this binary is suid, I’ve seen it on many systems so I believe that it is a common occurrence and not exploitable”. This is alluded to common-sense or experience by pen-testers (Happe and Cito, 2023b). GPT-4-turbo does not have this experience yet. The same behavior of testing all potential suid binaries, was seen while using the same vulnerable virtual machines with novice human penetration testers.
6.6 Threats to Validity
@@ 7 Experience and Guidance
  4. 4.
    Larger context sizes yielded better results but within our use-cases, context size usage oftentimes stagnated at approximately 20k tokens, indicating that massive context sizes might not be required for penetration-testing. Using in-context learning substantially increased costs while not significantly improving success rates.
  5. 5.
    Enumeration-based guidance was not as effective as high-level guidance, partially due to enumeration tools “staying in the box”. Only the most expressive models (e.g., gpt-4-turbo) were able to extract sufficient guidance from the enumeration tool’s output. Combining gpt-4-turbo for enumeration analysis with gpt-3.5-turbo for execution command generation yielded a cost-effective hybrid that was able to solve 40% of challenges.
  6. 6.
@@
8 Conclusion
There is both academic and industrial interest in integrating LLMs with penetration-testing. The efficient usage of LLMs depends on a firm understanding of their capabilities and strengths. To bolster this understanding, we have created an automated LLM escalation prototype and evaluated multiple LLMs. We gained insights into their capabilities, explored the impact of different history strategies, analyzed the quality of generated commands, and compared results with human hackers. We also released our created benchmark to foster further automation research.
@@ 8 Conclusion
The source code of hackingBuddyGPT can be found at https://github.com/ipa-lab/hackingBuddyGPT. The benchmark can be found at https://github.com/ipa-lab/benchmark-privesc-linux.
Conflict of Interest
The authors declare that they have no conflict of interest.